OpenStudy (ansleyfaithechols):

X 1 2 3 4 5 6 7 8 P(X) 1/8 1/8 1/8 1/8 1/8 1/8 1/8 1/8 What is the standard deviation of the random variable X, "the number that comes up on the die"? A. 3.745 B. 3.5 C. 1.708 D. 14.028 E. 2.291

OpenStudy (holsteremission):

The variance for the distribution is given by \[\mathbb E(X^2)-\mathbb E(X)^2=\sum_x x^2p(x)-\left(\sum_xxp(x)\right)^2\]and the standard deviation is the square root of the variance.